OpenStudy (lxelle):

A curve has the parametric equations x=1 - 2p and y = 2/p - 2. a) Find the equation of the noraml to the curve at the point Q(-1,0). b) The normal to the curve at the point Q cuts the curve again at the point R. Find the coordinates of point R.

OpenStudy (lxelle):

@perl @phi

OpenStudy (tkhunny):

Clarify y(p) = 2/p - 2 = \(\dfrac{2}{p} - 2\), as you have written it, or y(p) = \(\dfrac{2}{p-2}\).